Clinical Trials

Will King

Background

Causal Story and Data

Formalization Data Sources

Analysis

Econometric Model Results Total Effect Direct Effects

Conclusion

Appendices

Pharmaceutial competitors and their effect on clinical trial completion

Will King

Washington State University

william.f.king@wsu.edu

October 4, 2023

1 / 19 Clinical Trials

Will King

Background

Causal Story and Data

Formalization Data Sources

Analysis

Econometric Model Results Total Effect Direct Effects

Conclusion

Appendices

Introduction

2 / 19 Clinical Trials

Will King

Background

Causal Story and Data

Formalization Data Sources

Analysis

Econometric Model Results Total Effect Direct Effects

Conclusion Appendices

Overview I

1 Background
2 Causal Story and Data

Formalization Data Sources

3 Analysis

Econometric Model Results

Total Effect Direct Effects

4 Conclusion
5 Appendices

3 / 19 Clinical Trials

Will King Background

Causal Story and Data

Formalization Data Sources

Analysis

Econometric Model Results Total Effect Direct Effects

Conclusion Appendices

Clinical Trials and Drug develoment

The FDA requires clinical trials before approving new drug compounds Pre-registered design Updated regularly on clinicaltrials.gov Often requires an oversight board. Goal is to prove efficacy and safety of a compound/dosage/route. A new drug candidate (NDC) must complete 3 phases of clinical trials before approval. Phases are reviewed with FDA. Not all clinical trials are for new drugs.

4 / 19 Clinical Trials

Will King

Background

Causal Story and Data

Formalization Data Sources

Analysis

Econometric Model Results Total Effect Direct Effects

Conclusion

Appendices

Literature

Ekaterina Adams

5 / 19 Clinical Trials

Will King

Background

Causal Story and Data

Formalization Data Sources

Analysis

Econometric Model Results Total Effect Direct Effects

Conclusion Appendices

Questions of Interest

How do the competitors on the market affect clinical trial completion? How is this effect moderated by the enrollment of participants?

6 / 19 Clinical Trials

Will King

Background

Causal Story and Data

Formalization Data Sources

Analysis

Econometric Model Results Total Effect Direct Effects

Conclusion

Appendices

Audience Questions

What can I clarify?

7 / 19 Clinical Trials

Will King

Background

Causal Story and Data

Formalization Data Sources

Analysis

Econometric Model Results Total Effect Direct Effects

Conclusion Appendices

How do clinical trials proceed?

What does a complete trial look like. 1 Study sponsor comes up with design 2 Apply for NCT ID from ClinicalTrials.gov 3 Begin enrolling participants 4 Update ClinicalTrials.gov to recruit 5 Close Enrollment 6 Update ClinicalTrials.gov as not

recruiting* 7 Reach primary objectives 8 Update ClinicalTrials.gov as complete 9 Reach secondary objectives 10 Update ClinicalTrials.gov with more

information What does an incomplete trial look like? 1 Study sponsor comes up with design 2 Apply for NCT ID from ClinicalTrials.gov 3 Begin enrolling participants 4 Update ClinicalTrials.gov to advertise 5 Run into issues:

Safety Efficacy Profitability Feasiblity (enrollment, PI leaves, etc.)

6 Close Enrollment*

7 Decide to terminate clinical trial.

8 Update ClinicalTrials.gov as terminated.

8 / 19 Clinical Trials

Will King Background

Causal Story and Data

Formalization Data Sources

Analysis

Econometric Model Results Total Effect Direct Effects

Conclusion Appendices

ClinicalTrials.gov

Thus ClinicalTrials.gov becomes an (append only) repository of the “current” status of clincal trials. As it is designed to help faciltate enrollment in clinical trials, the record includes important information such as

drugs study arms conditions expected and final enrollment figures current status

ClinicalTrials.gov also reports the “history” from previous updates.

9 / 19

Will King

Background

Causal Story and Data

Formalization Data Sources

Analysis

Econometric Model Results Total Effect Direct Effects

Conclusion

Appendices

10 / 19 Clinical Trials

Will King

Background

Causal Story and Data

Formalization Data Sources

Analysis

Econometric Model Results Total Effect Direct Effects

Conclusion

Appendices

Questions?

11 / 19 Clinical Trials

Will King

Background

Causal Story and Data

Formalization Data Sources

Analysis

Econometric Model Results Total Effect Direct Effects

Conclusion

Appendices

12 / 19 Clinical Trials

Will King

Background

Causal Story and Data

Formalization Data Sources

Analysis

Econometric Model Results Total Effect Direct Effects

Conclusion

Appendices

13 / 19 Clinical Trials

Will King

Background

Causal Story and Data

Formalization Data Sources

Analysis

Econometric Model Results Total Effect Direct Effects

Conclusion

Appendices

14 / 19 Clinical Trials

Will King

Background

Causal Story and Data

Formalization Data Sources

Analysis

Econometric Model Results Total Effect Direct Effects

Conclusion

Appendices

15 / 19 Clinical Trials

Will King

Background

Causal Story and Data

Formalization Data Sources

Analysis

Econometric Model Results Total Effect Direct Effects

Conclusion

Appendices

16 / 19 Clinical Trials

Will King

Background

Causal Story and Data

Formalization Data Sources

Analysis

Econometric Model Results Total Effect Direct Effects

Conclusion

Appendices

17 / 19 Clinical Trials

Will King

Background

Causal Story and Data

Formalization Data Sources

Analysis

Econometric Model Results Total Effect Direct Effects

Conclusion

Appendices

Final Questions

Time is yours to ask any remaining questions.

18 / 19 Clinical Trials

Will King

Background

Causal Story and Data

Formalization Data Sources

Analysis

Econometric Model Results Total Effect Direct Effects

Conclusion Appendices

Equilibrium in the Money Market

Interest rates are the price of money, so we need to compare interest rates to the quantity of money demanded.

Figure: Money Market Equilibrium

19 / 19